OpenStudy (anonymous):

evaluate 3^2 + (8-2) times 4 - 6/3 A 30 B 58 C 28 D 31

OpenStudy (freckles):

\[3^2+(8-2) \cdot 4- \frac{6}{3}\] apply order of operations must do the operation inside the grouping symbol first

OpenStudy (freckles):

Grouping symbols examples of grouping symbols ( ), [ ] , | |

OpenStudy (freckles):

Do you see any grouping symbols?

OpenStudy (anonymous):

um nope

OpenStudy (freckles):

what about the ( )

OpenStudy (freckles):

I see 8-2 inside it

OpenStudy (anonymous):

yes the ( ) sign is there

OpenStudy (freckles):

is that not a grouping symbol?

OpenStudy (anonymous):

oh wait ya oops haha

OpenStudy (freckles):

anyways perform the operation(s) inside the grouping symbol first (8-2)=(what)

OpenStudy (freckles):

if you have 8 dollars and someone takes away 2 how many dollars do you have left?

OpenStudy (anonymous):

sorry i wasn't online for a moment and 6

OpenStudy (freckles):

\[3^2+(8-2) \cdot 4-\frac{6}{3} \\ 3^2+( 6 ) \cdot 4 -\frac{6}{3}\] We performed any operations inside grouping symbols first. ------------ Now we want to look at exponents. Do you see any exponents?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

So we must evaluate the part with exponent 3^2=? remember 3^2 means 3*3 not 3*2

OpenStudy (freckles):

\[3^2+(8-2) \cdot 4-\frac{6}{3} \\ 3^2+( 6 ) \cdot 4 -\frac{6}{3} \\ 9+(6) \cdot 4-\frac{6}{3}\] Second step way take care of the exponents. ----- Third, we want to perform multiplication/division as it occurs from left to right so 6(4)=? and 6/3=? --- Then last perform addtion/subtraction as it occurs left to right --- See if you can finish up and I will check your answer.